OpenStudy (roadjester):

Solve the matrix. 2 1 8 5 4 6 1 3 1

OpenStudy (anonymous):

There is no solution

OpenStudy (roadjester):

oops, my bad 2 1 8 1 4 2 5 4 6 * 2 1 1 1 3 1 1 0 1

OpenStudy (roadjester):

Forgot the second part. 'sweatdrops anime style.'

OpenStudy (anonymous):

the answer should be in the form of x1 x2 x3 Do you agree?

OpenStudy (anonymous):

2 1 8 1 4 2 5 4 6 * 2 1 1 = 1 3 1 1 0 1 (2*1)+ (1*4)+ (8*2) 22 (5*2)+ (4*1)+ (6*1) = 20 (1*1)+ (3*0)+ (1*1) 2

OpenStudy (roadjester):

Uh isn't it row*column? That's what I hate about a matrix. So confusing...

OpenStudy (roadjester):

heh, not the most conventional method, but my graphing calculator apparently can solve a matrix... I'll leave this up for a little while just in case people are viewing. I'll delete it later on.

OpenStudy (anonymous):

Am I wrong? I'm confused.

OpenStudy (roadjester):

Don't worry, I'm confused too. If anyone is well-versed in matrices, please help. I'm going to remove this question in a few hours and just ask my teacher/professor. Good thing it's a passing thing and we aren't going to go into too much detail.

OpenStudy (anonymous):

If A is an mXn matric and if B is an nXp matric then the product of AB is the mXp matrix so the answer should be a 3* 1